OpenStudy (anonymous):

find any points at which the function is discontinuous: y=(sqaure root of (x^4+1)/(1+sin^2x

OpenStudy (anonymous):

\[ f(x) = \frac{\sqrt{x^2+1}}{1+\sin^2 x} \] is continous on the real line.

OpenStudy (anonymous):

what does that mean? and where did the x^4 go inside the square root?

OpenStudy (anonymous):

sorry its meant to be a x^4 :D

OpenStudy (anonymous):

but that doesn't make a difference

OpenStudy (anonymous):

but that doesnt give me any points where the function could be discontinuous

OpenStudy (anonymous):

it means that no matter what real number you put in for x, that 1) there is not going to be a negative number below the radical, as x^4+1>0 2) there is not going to be a zero in the denominator because sin^2(x) > 0

OpenStudy (anonymous):

There are NO real numbers for which this is discontinuous!

OpenStudy (anonymous):

ok ok ok i gotcha. so how do i show my work for that problem? just write down those 2 resons you just gave me?

OpenStudy (anonymous):

yes and hit good answer